# Thermal Management Design and Key Technology Validation for PandaX Underground Experiment

### v1

@@ Abstract
The scale of liquid xenon experiments for rare events searching is expanding, which is planned even to fifty tons level. The detector and distillation tower require a reliable cooling source with large cooling power at liquid xenon temperature range. Pulse tube refrigerators and GM refrigerators, which were widely used in previous detectors, have the disadvantages of small cooling power, large space occupation, and non-standby mutuality, which become bottlenecks of the experiment scale expansion. In this study, an auto-cascade refrigerator with ethanol coolant is developed, and the heat transfer effect is improved by adopting the concentric shaft heat exchanger and after-pumping heat transfer scheme. The 2.5 kW stable cooling power is obtained at 155 K. Further, the feasibility and key technology of the centralized cooling system of 5 kw at 160 K is discussed. The study can simplify liquid xenon experimental auxiliary devices, which will be helpful for the PandaX-xT experiment scheme and its laboratory infrastructure design.
@@
journal: Energy
@@ 1 Introduction
Direct detection of dark matter is at the forefront of particle physics experiments, with liquid xenon detectors being the most promising technology. The scale of the detectors has developed from the initial 10kg to the current 10-ton level gradually, such as PandaX[1], XENON[2][3], LZ(LUX)[4]. The scale would reach fifty tons in the long-term plan [5],[6],[7]. Furthermore, the detection target has extended from dark matter to neutrino-less double beta decay events as well [6][8]. To reduce the background from cosmic rays, these experiments must operate in underground laboratories with a long construction period, high cost, and relatively limited space [9]. So, the sizes of the equipment underground, especially surrounding the detector, must be considered in the system design of PandaX-xT.
@@
To improve the detector’s sensitivity, the distillation tower for removing Rn and Kr is developed to operate online, and a cold electronic system might be installed inside the detector, increasing the cooling power requirements greatly. The operation temperature of the liquid xenon detector is about 178K. The pulse tube refrigerator (PTR) and GM refrigerators are utilized to supply the cooling power to the detector [10] and the distillation system generally [11]. Liquid nitrogen is used as the powerful auxiliary cooling source for xenon injection and recovery and in emergency conditions. However, the PTR and GM refrigerators are generally oriented to deep temperature refrigeration below 77K, which is expensive and has relatively small cooling power, which does not match the temperature requirements of the liquid xenon detector well. The cooling power allowance is required during design, and the cooling output is adjusted by heating or matching with the thermal load to achieve precise temperature control during operation [10]. In this case, different subsystems, such as the detector and distillation tower, can not share one PTR or GM refrigerator, resulting in the surplus cooling power not supporting each other mutually and the cooling power being wasted. In addition, liquid nitrogen supply is not as convenient as on the ground in China Jinping underground Laboratory (CJPL), so it is infeasible to obtain large cooling power using liquid nitrogen. In conclusion, using a liquid nitrogen-assisted refrigeration scheme based on PTR and GM refrigerators for small-scale liquid xenon experiments is feasible; however, its limited cooling power and low efficiency can not satisfy the experiment’s scale expansion requirements. Therefore, it is necessary to develop thermal management technology for the entire liquid xenon experimental system, adopt a centralized cooling scheme (CCS) and high cooling power refrigeration based on electric energy, improve the efficiency of cold energy utilization, reduce the space and human resource occupation of refrigeration equipment, improve the automation and reliability of the experimental system, and focus on the key task of particle detection.
@@
Section 2 summarizes the requirements of various cooling sources in the PandaX experiment. Section 3 illustrates the design, manufacturing, and performance testing results of the auto-cascade refrigerator (ACR). Section 4 discusses the design and feasibility analysis of the CCS. Section 5 discusses the precision temperature control cold head based on ACR (CHACR). We conclude in section 6.
@@ 2 Cooling Source Requirements of PandaX Experiment
In the PandaX experiments, the cooling requirements for a steady-state include the normal operation of the detector and distillation tower, with temperature stability of 178±0.1 K[10],[11]; The transient cooling requirements is pre-cooling of xenon gas and detectors, without temperature stability requirements [5],[12]. The xenon gas recovery stage before the detector ends operation requires room temperature ethanol as the heat source for liquid xenon vaporization [13]. The xenon storage system requires both cooling source and heat soure too[5],[12],[14]. Table 1 lists the main technical parameters of the refrigerators selected in the PandaX-4T experiment[10] and the PandaX-xT scheme [6], which need 1 kW and 3.4 kW cooling power, respectively.
@@
Experimental device Cooling power Refrigerator model Lowest Output power Cooling Power Electricity
@@
PT-90 (Cryomech, USA) 32 K 90 W @ 80 K 140 W 4.3 kW
PC-150 (JEC, Japan) 77 K 170 W @ 165 K 200 W 5.7 kW
PandaX-4T Distillation Tower 400 W AL-300 (Cryomech, USA) 25 K 320 W @80 K 500 W 7 kW
@@
PandaX-xT Cooling Bus 1300 W AL-600 (Cryomech, USA) 25 K 600 W @ 80 K 950 W 11.5 kW
@@
PandaX-xT Rn Distillation Tower 1300 W AL-600 (Cryomech, USA) 25 K 600 W @80 K 950 W 11.5 kW
@@ 3 Auto-Cascade Refrigeration (ACR)
The CCS requires a centralized high-power refrigeration system or cooling source; one scheme is to use liquid nitrogen as the primary cooling source and then control the temperature [15], which is feasible in ground experiments where liquid nitrogen supply is convenient. However, getting large quantities of liquid nitrogen in CJPL is not convenient, and there is no suitable equipment to produce liquid nitrogen at the CJPL. Considering refrigeration efficiency, using liquid nitrogen to liquify xenon gas is very inefficient due to the huge temperature difference. Electricity is the most convenient and reliable energy at the CJPL, so the electricity-based refrigerator is more suitable for PandaX experiments. The ACR can realize multistage overlapping utilizing a mixture of refrigerant and a compressor and obtain a low temperature below 160 K using an ethanol coolant[16], which facilitates the distribution and management of cooling capacity. The temperature difference of 18 K considers both ACR refrigeration efficiency and the feasibility of CHACR, corresponding to the PandaX 178 K operating temperature.
3.1 ACR Requirements and Design
@@
According to the discussion above, using two ACR units in operation with one as a backup can balance cost-effectiveness and reliability, so the single ACR design cooling power should be more than 2 kW at 160 K.
@@
The working principle of the ACR is shown in figure 1. The ethanol coolant is driven by gear pump A (Model NP1700) from the coolant vessel to the coaxial heat exchanger for cooling, which is also the refrigerator’s evaporator. After cooling, the coolant returns to the vessel or directly transfers to the external thermal load. The total volume of the coolant vessel is about 50 L, which is divided by a plate into two parts of 40 L and 10 L with temperature sensors A and B, respectively. There is a gap of about 1 cm between the dividing plate and the inner wall for coolant flow. A filter of 400 meshes is installed at the bottom of the vessel to protect gear pump A, which is driven by a servo motor to adjust the rotary speed continuously, and its rated flow rate is 30 L/min at 2000 r/min. There is an electric heater A between the gear pump A and the coaxial heat exchanger for ice melting. The upper vessel is connected to gear pump B, and the coolant can also be pumped out to the external thermal load. An electric heater B is installed at the output pipeline, which is used for heating mode. The ACR working principle is described in detail, combined with different operation modes as follows.
Figure 1: Design principle diagram of ACR
@@ 3.1.1 Pre-cooling Mode
It takes about 2 hours to cool the coolant from room temperature to 160 K in the coolant vessel. The ACR can start in this mode for pre-cooling to save time. The rated temperature of the ACR is 160K, so the compressor load is large when the temperature of the coolant flowing through the evaporator is much higher than the rated temperature. The deviation from the rated condition shortens the ACR’s life. To solve the problem, it is designed to cool the coolant in the lower secondary refrigerant vessel first, and then the coolant in the upper vessel is cooled gradually.
For example, the ACR started from room temperature with the default operation parameters. In the pre-cooling mode, gear pump A delays starting for a few seconds, and the coolant temperature in the lower vessel determines the flow rate. The flow rate is 5 L/min, 10 L/min, 20 L/min and 30 L/min corresponding to the temperature higher than 273 K, from 223 K to 273 K, from 173 K to 223 K, and lower than 173K, respectively. If the coolant temperature in the lower vessel is above 173 K, the coolant cooled by the evaporator enters the lower vessel via valve B to be further cooled to below 173 K soon because of the smaller volume with the smaller heat capacity of the lower vessel. Then valve A opens, and the cryogenic coolant enters the upper vessel to gradually cool down its internal coolant. Meanwhile, the coolant with a higher temperature in the upper vessel flows into the lower vessel via the gap around the dividing plate and mixes with the low-temperature coolant, causing the temperature to increase. When the temperature rises above 173 K, valve A closes while valve B opens to cool down the lower vessel’s coolant preferentially. During the pre-cooling mode, the status of valves A and B switches several times, and the temperature of the upper vessel drops to 173K gradually. Afterward, the stable cooling process starts with valve A opening and valve B closing until the coolant cools to the setting temperature.
@@ 3.1.2 Indirect output Mode
This operation mode outputs the cooling coolant to cool down the thermal load, which can be considered an upgraded pre-cooling mode. After the indirect output mode starts, the gear pumps A and B are switched on simultaneously, and the cooling process is consistent with the pre-cooling mode. Valves D and F open while pump B starts, and the thermal load temperature is measured by temperature sensor E after 1 minute. Suppose the temperature difference between sensors A and E is large, more than 30K, for instance. In that case, valve E opens intermittently to reduce the thermal stress, and the temperature difference determines the opening duty rate. The cold coolant in the upper vessel is mixed with the reflux coolant and compressed to the thermal load by gear pump B for cooling. The valve E opens, and valve F closes when the temperature difference between sensors A and E decreases to the setting value, such as 10K, to obtain better heat exchange between the cold coolant in the upper vessel and the thermal load. Thus, the stable cooling output from the vessel is realized for the thermal load. At the same time, the refrigerator also cools the coolant in the vessel, and valve C opens when the temperatures of the upper vessel and the reflux flow drop to 173 K. Valves A, B, and D, as well as gear pump B, are closed in order to cool down the thermal load directly with the coldest coolant from the evaporator, improving cooling efficiency. In this mode, the stability during the thermal load cooling process is ensured, and the negative influence of the large thermal load with high temperature on the refrigerator life is avoided too.
@@ 3.1.3 Direct output Mode
In this mode, the gear pump A starts, and valves A, B, D, F, and H close with valves C and E opening, then the compressor starts cooling after a delay of 1 minute, regardless of the current temperatures of the coolant and thermal load. In this mode, the coldest coolant directly flows to the thermal load to cool it down faster than the indirect out mode. This mode is suitable for rapid cooling requirements.
@@ 3.1.4 Deicing Mode
The ethanol in the evaporator has the risk of being solidified when the temperature gets close to or lower than its freezing point, for the designed operating temperature of the refrigerator is low. Once the ethanol solidifies in the evaporator, the heat transfer deteriorates, with the flow resistance increasing and the flow rate decreasing, further worsening the heat transfer. The heat exchanger enters a positive feedback state, rapidly losing its cooling capacity. Two pressure sensors are installed before and after the evaporator for monitoring, and the evaporator is considered to be frozen when the pressure difference exceeds 4 bar, then the ACR enters the deicing mode automatically. When the ACR operates in the indirect output mode, the gear pump B, valves D, and E open, while valves C and F close, and the thermal payload is still cooling by the coolant in the upper vessel. When the ACR operates in pre-cooling mode, the valves A and B close, and the valve H opens. After the deicing mode starts, valves A and B close while valve H opens, and the electric heater of 5kw is switched on, then the flow rate of gear pump A reduces to 2 L/min. The temperature of the heated ethanol is higher than 190 K with a volume of about 5 L in the evaporator. Thus, all the solid ethanol can be melted in 3 minutes at the deicing mode, and the heat exchange function of the evaporator was recovered. The deicing mode exits after 3 minutes automatically, then back to normal operation.

### v2

@@ Abstract
The scale of liquid xenon experiments aimed at searching for rare events is expanding, with plans to reach up to fifty tons. These detectors and the accompanying distillation towers require a reliable cooling source that can deliver large cooling power in the liquid xenon temperature range. Previously utilized pulse tube refrigerators and GM refrigerators have the disadvantages including small cooling power, large space occupation, and non-standby mutuality, which become significant challenges to scaling up experiments. In this study, an auto-cascade refrigerator was developed using ethanol as a coolant, with improved heat transfer effect achieved by adopting a concentric shaft heat exchanger and an after-pumping heat transfer scheme. This system provides a stable cooling power of 2.5 kW at 155 K. Further, the feasibility and key technology of the centralized cooling system of 5 kW at 160 K is discussed. This study promises to simplify liquid xenon experimental auxiliary devices, thereby benefiting the PandaX-xT experiment scheme.
@@
journal: International Journal of Refrigeration
@@ 1 Introduction
The direct detection of dark matter is a leading particle physics experiment, with liquid xenon detectors merging as the most promising technology. The scale of the detectors has developed from the initial 10 kg to the current 10-ton level gradually, such as PandaX[1], XENON[2][3], LZ(LUX)[4]. The long-term plan aims to increase this scale to fifty tons [5],[6],[7]. Furthermore, the detection target has extended from dark matter to neutrino-less double beta decay events [6][8]. To reduce the background from cosmic rays, these experiments must operate in underground laboratories with a long construction period, high cost, and relatively limited space [9]. Consequently, the dimensions of the equipment underground, especially around the detector, must be carefully considered in the system design of PandaX-xT.
@@
To improve the sensitivity of the detector, an online distillation tower for removing Rn and Kr has been developed, and a plan to integrate a cold electronic system within the detector has been made, which significantly increases the cooling power requirements. The operational temperature of the liquid xenon detector is approximately 178 K. Liquid nitrogen is used as the powerful auxiliary cooling source for xenon injection and recovery in emergency conditions. However, pulse tube refrigerators (PTR) and GM refrigerators are commonly designed for deep temperature refrigeration below 77 K. These system is expensive with relatively small cooling power (usually below 1000 W), which makes them inadequate for the temperature requirements of the liquid xenon detector. During the design phase, it is essential to cooling power allowance, as the cooling output needs to be adjusted by heating or matching with the thermal load to achieve precise temperature control during operation [10]. Consequently, different subsystems, such as the detector and the distillation tower, could not share single PTR or GM refrigerator. This leads to excessive cooling power not supporting each other mutually and the cooling power being wasted. Furthermore, liquid nitrogen is not as readily accessible at the China Jinping Underground Laboratory (CJPL), making it impractical to rely on liquid nitrogen for substantial cooling power. In conclusion, using a liquid nitrogen-assisted refrigeration scheme using PTR and GM refrigerators for small-scale liquid xenon experiments is feasible. However, inherent limitations in cooling power and efficiency are insufficient to meet the demands of scaling up the experiments. Therefore, it is essential to develop the thermal management technology for the entire liquid xenon experimental setup by adopting a centralized cooling scheme (CCS) and high cooling power refrigeration based on electric energy. This approach aims to improve the cold energy utilization efficiency, reduce the space and labor occupation of refrigeration equipment, improve the automation and reliability of the experimental system, and focus on the key task of particle detection.
@@
Section 2 outlines the cooling source requirements for the PandaX experiment. Section 3 presents the design, manufacturing, and performance testing results of the auto-cascade refrigerator (ACR). Section 4 discusses the design and feasibility analysis of the CCS. Section 5 discusses the precision temperature control cold head based on ACR (CHACR). The discussion is concluded in section 6.
@@ 2 Cooling Source Requirements of PandaX Experiment
In the PandaX experiments, the cooling requirements for a steady-state include the normal operation of the detector and the distillation tower, necessitating a temperature stability of 178±0.1 K[10],[11]; For transient cooling requirements, pre-cooling of xenon gas and detectors is essential, though temperature stability is not a requirement in this case [5],[12]. During the xenon gas recovery stage before the detector ends operation requires room temperature ethanol as the heat source for liquid xenon vaporization [13]. Additionally, the xenon storage system requires both a cooling source and heat source [5],[12],[14]. Table 1 outlines the key technical parameters of the refrigerators selected for the PandaX-4T experiment[10] and the PandaX-xT scheme [6], which need 1 kW and 3.4 kW cooling power, respectively.
@@
Experimental device Cooling power Refrigerator type and model Lowest Output power Cooling Power Electricity
@@
Pulse tube refrigerator PT-90 (Cryomech, USA) 32 K 90 W @ 80 K 140 W 4.3 kW
GM refrigerator PC-150 (JEC, Japan) 77 K 170 W @ 165 K 200 W 5.7 kW
PandaX-4T Distillation Tower 400 W GM refrigerator AL-300 (Cryomech, USA) 25 K 320 W @80 K 500 W 7 kW
@@
PandaX-xT Cooling Bus 1300 W GM refrigerator AL-600 (Cryomech, USA) 25 K 600 W @ 80 K 950 W 11.5 kW
@@
PandaX-xT Rn Distillation Tower 1300 W GM refrigerator AL-600 (Cryomech, USA) 25 K 600 W @80 K 950 W 11.5 kW
@@ 3 Auto-Cascade Refrigeration (ACR)
The CCS requires a centralized high-power refrigeration system or cooling source. One proposed approach is to use liquid nitrogen as the primary cooling source and then control the temperature [15]. This is feasible in ground experiments where liquid nitrogen is readily available. However, sourcing large quantities of liquid nitrogen at the CJPL poses a logistical challenge, and there is now no suitable equipment for liquid nitrogen production on site. When considering refrigeration efficiency, using liquid nitrogen to liquefy xenon gas is highly inefficient due to the significant temperature difference involved. Given that electricity is the most accessible and reliable energy at the CJPL, an electricity-based refrigerator is a more suitable option for PandaX experiments. The ACR could achieve multistage overlapping by utilizing a mixed refrigerant and a single compressor. This enables it to reach a temperature below 160 K with an ethanol coolant[16]. This facilitates effective distribution and management of cooling capacity. The 18 K temperature difference accounts for both ACR refrigeration efficiency and the feasibility of CHACR, aligning with the PandaX operating temperature of 178 K.
3.1 ACR Prinple and Design
@@
Based on the above discussion, using two ACR units in operation with one as a backup can balance cost-effectiveness and reliability. Consequently, the cooling power of a single ACR design should exceed 2 kW at 160 K.
The ACR unit employs a mixed refrigerant and a single compressor to facilitate a multi-stage cascade, achieving an evaporation temperature of approximately 150 K. The operation principle is illustrated in figure 1. Initially, the mixed refrigerant is compressed into a high-temperature and high-pressure state by the compressor. It then undergoes cooling through a water-cooled condenser and a regenerative heat exchanger, resulting in the liquefaction of some high boiling point components. In contrast, the low boiling point components remain gaseous. Then the mixed refrigerant of the gas-liquid phase enters the first stage gas-liquid separator, and the liquid part enters the first stage intermediate heat exchanger after being depressurized and cooled via a throttle valve. This process further cools the low boiling point gas phase components from the gas-liquid separator. The two-phase mixed refrigerant then flows into the subsequent gas-liquid separator. This sequence is repeated until the component with the lowest boiling point is completely liquefied in the fourth stage intermediate heat exchanger. Subsequently, it was throttled and evaporated, to achieve a temperature of 150 K. Afterwards, it entered the final stage evaporator to cool the flowing coolant[16] . The gaseous refrigerant exiting the intermediate heat exchanger at each is compressed and pressurized by the compressor after flowing through the heat exchanger, thereby completing the refrigeration cycle. The ACR could achieve effective refrigeration output below 160 K, which greatly enhances distribution management and simplifies the cooling system. Compared with the detector operating temperature of 178 K, the 18 K temperature difference margin balances the refrigeration efficiency and realizes the precision temperature control modules based on heat conduction.
Figure 1: Working principle of ACR
@@
The coolant circulation principle of the ACR is illustrated in figure 2. The ethanol coolant is driven by gear pump A (Model NP1700) from the coolant vessel to the coaxial heat exchanger, which function as the refrigerator’s evaporator for cooling. After cooling process, the coolant either returns to the vessel or is directly transferred to the external thermal load. The total volume of the coolant vessel is about 50 L, which is divided by a plate into two parts of 40 L and 10 L with temperature sensors A and B, respectively. There is a gap of about 1 cm between the dividing plate and the inner wall for coolant flow. A filter of 400 meshes is installed at the bottom of the vessel to protect gear pump A. Gear pump A is driven by a servo motor, facilitating continuous adjustment of the rotary speed, boasting a flow rate of 30 L/min at 2000 r/min. An electric heater A is positioned between gear pump A and the coaxial heat exchanger to aid in ice melting. The upper vessel is connected to gear pump B, enabling the coolant to be pumped out to the external thermal load. Furthermore, an electric heater B is installed at the output pipeline, utilized during heating mode. The ACR working principle is described in detail, combined with various operational modes as follows.
Figure 2: Coolant circulation diagram of ACR
@@ 3.1.1 Pre-cooling Mode
Cooling the coolant from room temperature to 160 K in the coolant vessel takes approximately 2 h. To expedite this process, the ACR can start pre-cooling in this mode. The ACR is rated for a temperature of 160 K, resulting in a significant compressor load when the coolant flowing through the evaporator is at a much higher temperature. The deviation from the rated condition shortens the lifespan of ACR. To address this issue, the design incorporates a two-step cooling process: first, the coolant in the lower vessel is cooled, followed by a gradual cooling of the coolant in the upper vessel.
For instance, the ACR begins at room temperature with the default operation parameters. During the pre-cooling mode, gear pump A delays its activation for a few seconds, and the coolant temperature in the lower vessel determines the flow rate. The flow rates are 5 L/min, 10 L/min, 20 L/min and 30 L/min corresponding to the temperature higher than 273 K, from 223 K to 273 K, from 173 K to 223 K, and lower than 173 K, respectively. If the coolant temperature in the lower vessel is above 173 K, the coolant cooled by the evaporator enters the lower vessel via valve B to be further cooled to below 173 K soon because of the smaller volume with the smaller heat capacity of the lower vessel. Subsequently, valve A opens, allowing the cryogenic coolant to flow into the upper vessel, gradually lowering the temperature of its internal coolant. Simultaneously, the coolant with a higher temperature in the upper vessel flows into the lower vessel via the gap around the dividing plate and mixes with the low-temperature coolant, causing the temperature to increase. Once the temperature exceeds 173 K, valve A closes while valve B opens to cool down the lower vessel’s coolant preferentially. During the pre-cooling mode, the status of valves A and B switches several times, and the temperature of the upper vessel drops to 173 K gradually. Afterward, the stable cooling process starts with valve A opening and valve B closing until the coolant reaches the desired setting temperature.
@@ 3.1.2 Indirect output Mode
This operation mode delivers coolant to effectively cool down the thermal load, functioning as an upgraded pre-cooling mode. Once the indirect output mode starts, the gear pumps A and B are switched on simultaneously, and the cooling process is consistent with the pre-cooling mode. Valves D and F open while pump B starts, and the thermal load temperature is measured by temperature sensor E after 1 min. Suppose the temperature difference between sensors A and E exceeds 30 K, valve E opens intermittently to reduce the thermal stress, and the temperature difference determines the opening duty rate. The cold coolant in the upper vessel is then mixed with the reflux coolant and delivered to the thermal load by gear pump B for cooling purposes. When the temperature difference between sensors A and E decreases to the predetermined value, such as 10 K, valve E opens while valve F closes. This adjustment enhances heat exchange between the cold coolant in the upper vessel and the thermal load, resulting in a stable cooling output for the thermal load. Simultaneously, the refrigerator continues to cool the coolant in the vessel, and valve C opens when the temperatures of the upper vessel and the reflux flow drop to 173 K. Valves A, B, and D, and gear pump B, are closed to cool down the thermal load directly with the coldest coolant from the evaporator, improving cooling efficiency. In this mode, stability during the thermal load cooling process is maintained, preventing any adverse effects of the high temperature thermal load on the lifespan of the refrigerator.
@@ 3.1.3 Direct output Mode
In this mode, gear pump A is activated, causing valves A, B, D, F, and H to close while valves C and E open. Following a 1 min delay, the compressor begins the cooling process, regardless of the current temperatures of the coolant and conditions of the thermal load. This mode allows the coldest coolant to flow directly to the thermal load facilitating faster cooling than the indirect mode. This mode is specifically designed for situations requiring rapid cooling.
@@ 3.1.4 Deicing Mode
The ethanol in the evaporator has the risk of being solidified when the temperature gets close to or lower than its freezing point, for the designed operating temperature of the refrigerator is low. Once the ethanol solidifies in the evaporator, the heat transfer deteriorates, with the flow resistance increasing and the flow rate decreasing, further worsening the heat transfer. The heat exchanger enters a positive feedback state, rapidly losing its cooling capacity. Two pressure sensors are installed before and after the evaporator for monitoring, and the evaporator is considered frozen when the pressure difference exceeds 4 bar. Subsequently, the ACR enters the deicing mode automatically. When the ACR operates in the indirect output mode, the gear pump B, valves D and E open, while valves C and F close, and the thermal payload is still cooling by the coolant in the upper vessel. When the ACR operates in pre-cooling mode, the valves A and B close, and the valve H opens. After the deicing mode starts, valves A and B close while valve H opens, and the electric heater of 5 kW is switched on, then the flow rate of gear pump A reduces to 2 L/min. The temperature of the heated ethanol rises above 190 K with a volume of approximately 5 L in the evaporator. This is sufficient to melt all solid ethanol within 3 min in deicing mode, thereby restoring the heat exchange function of the evaporator. The deicing mode automatically exits after 3 min, returning the system back to normal operation.